Salaries for teachers in a particular elementary school district are normally distributed with a mean of $44,000 and a standard deviation of $6,500. We randomly survey ten teachers from that district.

  1. Find the 90th percentile for an individual teacher’s salary.
  2. Find the 90th percentile for the average teacher’s salary.
Answer #1

1)

mean μ= 44000.000
standard deviation σ= 6500.000
for 90th percentile critical value of z= 1.28
therefore corresponding value=mean+z*std deviation=44000+1.28*6500= $ 46773.33

2)

mean μ= 44000.000
sample size       =n= 10
std error=σ=σ/√n= 2055.4805
for 90th percentile critical value of z= 1.28
therefore corresponding value=mean+z*std deviation=44000+1.28*2055.4805= $ 46631.02